# Nonlinear Causal Discovery through a Sequential Edge Orientation Approach

### v1

@@ Abstract
Recent advances have established the identifiability of a directed acyclic graph (DAG) under additive noise models (ANMs), spurring the development of various causal discovery methods. However, most existing methods make restrictive model assumptions, rely heavily on general independence tests, or require substantial computational time. To address these limitations, we propose a sequential procedure to orient undirected edges in a completed partial DAG (CPDAG), representing an equivalence class of DAGs, by leveraging the pairwise additive noise model (PANM) to identify their causal directions. We prove that this procedure can recover the true causal DAG assuming a restricted ANM. Building on this result, we develop a novel constraint-based algorithm for learning causal DAGs under nonlinear ANMs. Given an estimated CPDAG, we develop a ranking procedure that sorts undirected edges by their adherence to the PANM, which defines an evaluation order of the edges. To determine the edge direction, we devise a statistical test that compares the log-likelihood values, evaluated with respect to the competing directions, of a sub-graph comprising just the candidate nodes and their identified parents in the partial DAG. We further establish the structural learning consistency of our algorithm in the large-sample limit. Extensive experiments on synthetic and real-world datasets demonstrate that our method is computationally efficient, robust to model misspecification, and consistently outperforms many existing nonlinear DAG learning methods.
@@ 1 Introduction
Structural causal models (Pearl, 2000) represent the causal relations amongst a set of variables using a directed acyclic graph (DAG), while the underlying data generating process is described by a set of structural equation models (SEMs). In practice, the true DAG is often unknown or difficult to construct due to limited domain knowledge. Consequently, a wide range of causal discovery methods have been developed to learn the underlying DAG or its equivalence class from observational data (Glymour et al., 2019; Vowels et al., 2022).
@@
where $\text{PA}_{i}$ denotes the parent set of $X_{i}$. A common assumption is that each $f_{i}(\cdot)$ is a linear SEM with additive Gaussian noise. Although analytically convenient, this assumption is not only overly simplistic, but also limits algorithms to learning a completed partially direct acyclic graph (CPDAG), which encodes a set of Markov equivalent DAGs sharing the same conditional independence relations (Chickering, 2002), rather than the exact true DAG. Nevertheless, previous works such as Hoyer et al. (2008) and Shimizu et al. (2006) have demonstrated that the true DAG is identifiable under nonlinear functions and/or non-Gaussian noise terms in the SEM. These assumptions break the symmetry in the bivariate distribution of two nodes, enabling the identification of causal directions (Zhang and Hyvärinen, 2016).
@@ 1.1 Relevant Work
In this paper, we focus on learning causal DAGs from nonlinear data. Prior works, such as the additive noise model by Hoyer et al. (2008) and post-nonlinear model by Zhang and Hyvarinen (2012), have established the identifiability of the true DAG under particular assumptions on the function class $f_{i}(\cdot)$ in the general SEM. New causal discovery algorithms have then ensued from these identifiability results. In the domain of constraint-based methods, kernel-based tests have garnered considerable attention. The kernel-based conditional independence (KCI) test, a notable early example proposed by Zhang et al. (2011), captures nonlinear relationships by mapping random variables to reproducing kernel Hilbert spaces using kernel methods. Building on this, the algorithm RESIT (Peters et al., 2014) utilizes a kernel-based statistic to recursively identify sink nodes and infer the topological ordering of a DAG. Another line of work employs regression-based methods to detect nonlinearity. For instance, the nonlinear invariant causal prediction (ICP) framework (Heinze-Deml et al., 2018) fits regression models to data collected in different environments, then tests for differences in residual distributions across environments after interventions to detect associations. Improving upon an estimated CPDAG, Wang and Zhou (2021) propose a statistical test that orients the undirected edges by comparing the goodness of fit of models corresponding to the two possible directions. With minimal assumptions on the regression model, this approach offers flexibility in modeling and greater applicability to real-world data.
@@
However, these constraint-based methods also face practical limitations that hinder their widespread adoption. A significant drawback of kernel-based approaches is the computational cost. As its runtime scales quadratically with sample size, the KCI test is computationally expensive to employ in constraint-based algorithms, and especially inefficient for learning larger DAGs (Strobl et al., 2019). Additionally, the performance of the KCI test heavily depends on the chosen kernel function, which is often problem-specific and difficult to tune. The nonlinear ICP requires a sufficient number of observed environments to detect causal relations (Rosenfeld et al., 2021) and thus cannot be applied to data generated under a classical i.i.d. setting. Moreover, the accuracy of regression-based approaches is contingent on the quality of the estimated model, which requires strong domain knowledge for model selection and large sample sizes to accurately approximate the true SEM (Shah and Peters, 2020). Model misspecification can lead to violations of key assumptions, ultimately inflating false positive rates in independence tests (Li and Fan, 2020). More broadly, recent efforts tend to focus on developing independence tests- without fully developing them into scalable, full-fledged causal discovery algorithms (Hasan et al., 2023)- or approximating key statistical quantities to improve test efficiency, with less emphasis on using these metrics to infer causal relations.
@@
Score-based methods for nonlinear learning have also seen significant development in recent years. The algorithm CAM (Bühlmann et al., 2014), for instance, assumes an additive noise model with Gaussian errors and maximizes the joint log-likelihood function to learn a DAG. More recently, several works have reformulated the structural learning task as a continuous-optimization problem by devising an algebraic characterization of DAGs (Zheng et al., 2018). Prominent examples include NOTEARS (Zheng et al., 2020), DAG-GNN (Yu et al., 2019), and DAGMA (Bello et al., 2022), which incorporate deep learning techniques to enhance the flexibility of SEM estimation. Another line of work, exemplified by the SCORE algorithm (Rolland et al., 2022), iteratively identifies leaf nodes in a DAG using the Jacobian of the score function. However, score-based methods generally heavily depend on model assumptions, require intensive computational time, and are not guaranteed to find a globally optimal structure. Some of the above deep learning-based methods, in particular, perform poorly when the data is standardized, as standardization erases causal order information from the marginal variance when minimizing the least squares objective function (Reisach et al., 2021).
@@ 1.2 Contribution of This Work
In lieu of inferring the causal order of nodes, we introduce a novel identifiability criterion based on the pairwise additive noise model (PANM). This criterion determines whether a given undirected edge in a partially directed acyclic graph (PDAG) can be correctly oriented. Edges that fulfill this criterion are ensured to be correctly oriented in the large-sample limit, without inducing errors in subsequent orientations. We establish theoretical results proving that, at the population level, the algorithm consistently recovers the true DAG from a CPDAG. Leveraging this result, we devise a sequential algorithm to identify undirected edges following the PANM criterion and infer their causal directions. As such, the algorithm effectively learns the DAG at a local scale by evaluating edges individually without necessitating evaluation of all nodes or the entire graph. To determine the orientation of each edge, SNOE employs a likelihood-ratio test that compares the bivariate conditional probability distributions over the sub-DAG on both nodes and their learned parent sets under the competing directions. Contrary to general independence tests, the likelihood ratio test returns a definitive decision regarding the causal direction, thereby bridging the gap between the task of detecting nonlinear conditional independence relations and the structural learning problem. Empirical results further show that the test is robust to violations of model misspecification and yields accurate results across different functional settings.
@@
  • An algorithm that is guaranteed at the population level to identify the true DAG from its Markov equivalence class by orienting undirected edges satisfying the PANM criterion in a sequential manner;
@@
The paper is structured as follows. Section 2 introduces the fundamental concepts and assumptions. In Section 3, we present the central notion of our work- the pairwise additive noise model, followed by the sequential edge orientation procedure and the finite-sample version of the algorithm. In Section 4, we discuss the two key components of the edge orientation step: the ranking procedure and the orientation test. Then, we establish the structural learning consistency of the algorithm in Section 5. Section 6 presents the performance of our method against competing methods on both simulated and real-world data, with a detailed analysis of the algorithm’s intermediate results. Last, we summarize our work and outline directions for future research in Section 7.
@@ 2.1 Directed Acyclic Graph
A graph $\mathcal{G}=(V,E)$ consists of a set of vertices $V=[p]:=\{1,\ldots,p\}$ and a set of edges $E\subseteq V\times V$. For a pair of distinct nodes $i,j\in V$, a directed edge $i\rightarrow j\in E$ indicates that node $i$ is a parent of node $j$ and $j$ is a child of $i$. The parent set of node $i$ is denoted by $\text{pa}_{\mathcal{G}}(i)$ and the child set denoted by $\text{ch}_{\mathcal{G}}(i)$. In contrast, there may exist an undirected edge $i-j\in E$ in $\mathcal{G}$, where $i$ is called a neighbor of $j$, i.e. $i\in\text{ne}_{\mathcal{G}}(j)$, and vice versa. A directed acyclic graph (DAG) $\mathcal{G}$ consists only of directed edges and does not admit any directed cycles. A related graphical structure is the partially directed acyclic graph (PDAG), which contains both directed and undirected edges but does not contain any cycles in its directed subgraphs.
@@
To infer the structure of the graph $\mathcal{G}$ from the observed data, we require the causal sufficiency and faithfulness assumptions to hold (Pearl, 2000). Causal sufficiency is satisfied when all common causes of any distinct pair of nodes are observed. In other words, there are no unobserved or latent confounder under this assumption. Suppose $A,B,C\subset[p]$ are disjoint subsets. A DAG $\mathcal{G}$ and the probability distribution $P$ are faithful to one another if the conditional independence relations in $P$ have a one-to-one correspondence with the d-separation relations in $\mathcal{G}$: $A\perp\!\!\!\perp B|C\text{ in }P\Longleftrightarrow C$ d-separates $A,B$ in $\mathcal{G}$. Faithfulness also implies the causal minimality condition. The pair $(\mathcal{G},P)$ satisfies the causal minimality condition if $P$ is not Markov to any proper subgraph of $\mathcal{G}$ over the vertex set $V$ (Spirtes et al., 2000).
@@ 2.2 Markov Equivalence Class
Two DAGs are Markov equivalent if and only if they have identical skeletons and v-structures, which are ordered triplets of nodes $i,j,k$ oriented as $i\rightarrow k\leftarrow j$ with no edge between $i,j$ (Verma and Pearl, 1990). Markov equivalent DAGs encode the same set of d-separations and form an equivalence class. Without further restrictions on the function classes in the SEM (1), equivalent DAGs cannot be distinguished through observational data; hence we can only learn their equivalence class. The equivalence class is represented by a completed partially directed acyclic graph (CPDAG), a partial DAG with specific structural properties (Andersson et al., 1997). Every directed edge is compelled, or strongly protected, and every undirected edge is reversible in the CPDAG.
Since the CPDAG is also a chain graph, we can partition the vertex set $V$ into chain components as $V=V_{1}\cup...\cup V_{T}$ (Lauritzen, 1996). Edges between nodes within the same chain component are undirected, whereas edges between nodes in different chain components $V_{s},V_{t}$ are directed. The CPDAG $\mathcal{E}$ is typically obtained by first identifying v-structures in the skeleton of the DAG $\mathcal{G}$, and then applying Meek’s rules, a set of four rules that orient edges based on graphical patterns (Meek, 1995). Meek’s rules help to identify additional directed edges in the graph without introducing new v-structures. A maximally oriented partial directed acyclic graph (maximally oriented PDAG) is a PDAG for which no edges can be further oriented by Meek’s rules. As an example, a CPDAG $\mathcal{E}$ is a maximally oriented PDAG, or a maximal PDAG for short.
A consistent extension of a PDAG $\mathcal{G}$ is a DAG, $\widetilde{\mathcal{G}}$, obtained by orienting all undirected edges in $\mathcal{G}$ without introducing new v-structures. Therefore, $\widetilde{\mathcal{G}}$ has the same skeleton, the same orientations of all directed edges in $\mathcal{G}$, and the same v-structures as $\mathcal{G}$ (Dor and Tarsi, 1992). While not all PDAGs can be extended to a DAG, a CPDAG $\mathcal{E}$ is extendable since each DAG in the equivalence class represented by $\mathcal{E}$ is a consistent extension of $\mathcal{E}$. A DAG may be obtained from a CPDAG by iteratively making edge orientations without introducing new v-structures and applying Meek’s rules, while preserving the compelled edges (Wienöbst et al., 2021).

### v2

@@ Abstract
Recent advances have established the identifiability of a directed acyclic graph (DAG) under additive noise models (ANMs), spurring the development of various causal discovery methods. However, most existing methods make restrictive model assumptions, rely heavily on general independence tests, or require substantial computation. To address these limitations, we propose a sequential procedure to orient undirected edges in a completed partial DAG (CPDAG), representing an equivalence class of DAGs, by leveraging the pairwise additive noise model (PANM) to identify their causal directions. We prove that this procedure can recover the true causal DAG assuming a restricted ANM. Building on this result, we develop a novel constraint-based algorithm for learning causal DAGs under nonlinear ANMs. Given an estimated CPDAG, we develop a ranking procedure that sorts undirected edges by their adherence to the PANM, which defines an evaluation order of the edges. To determine the edge direction, we devise a statistical test that compares the log-likelihood values, evaluated with respect to the competing directions, of a sub-graph comprising just the candidate nodes and their identified parents in the partial DAG. We further establish the structural learning consistency of our algorithm in the large-sample limit. Extensive experiments on synthetic and real-world data sets demonstrate that our method is computationally efficient, robust to model misspecification, and consistently outperforms many existing nonlinear DAG learning methods.
@@ 1 Introduction
Structural causal models (pearl@2000) represent the causal relations amongst a set of variables using a directed acyclic graph (DAG), while the underlying data generating process is described by a set of structural equation models (SEMs). In practice, the true DAG is often unknown or difficult to construct due to limited domain knowledge. Consequently, a wide range of causal discovery methods have been developed to learn the underlying DAG or its equivalence class from observational data (glymour2019review; vowels2022d).
@@
where $\text{PA}_{i}$ denotes the parent set of $X_{i}$. A common assumption is that each $f_{i}(\cdot)$ is a linear SEM with additive Gaussian noise. Although analytically convenient, this assumption is not only overly simplistic, but also limits algorithms to learning a completed partially direct acyclic graph (CPDAG), which encodes a set of Markov equivalent DAGs sharing the same conditional independence relations, rather than the exact true DAG (chickering2002learning). Nevertheless, previous research has demonstrated conditions enabling the identifiability of the true DAG from observational data. Initial works proved bivariate identifiability under nonlinear functions and/or non-Gaussian noise terms in the SEM, as shown in hoyer2008nonlinear and shimizu2006linear, respectively. These assumptions break the symmetry in the bivariate distribution of two nodes, enabling the identification of causal directions (zhang2016nonlinear).
@@ 1.1 Relevant Work
In this paper, we focus on learning causal DAGs from nonlinear data. Prior works, such as the additive noise model (ANM) by hoyer2008nonlinear and post-nonlinear model by zhang2009pnl_identifiability, have established the identifiability of the true DAG under particular assumptions on the function class $f_{i}(\cdot)$ in the general SEM. peters2011IFMOC proved identifiability for the general SEM (1) by defining the concept of an identifiable functional model class (IFMOC). Specifically, the authors present a criterion on $\{X_{i},f_{i}(\text{PA}_{i}),\varepsilon_{i}\}$ for DAG identifiability, hence generalizing previous works focusing on specific models only. They also developed an algorithm to find all DAGs that satisfy their identifiability condition through iteratively testing for independence between residuals and parents.
@@
New causal discovery algorithms have then ensued from these identifiability results. In the domain of constraint-based methods, kernel-based tests have garnered considerable attention. The kernel-based conditional independence (KCI) test, a notable early example proposed by kci_test, captures nonlinear relationships by mapping random variables to reproducing kernel Hilbert spaces using kernel methods. Building on this, the algorithm RESIT (resit_anm) utilizes a kernel-based statistic to recursively identify sink nodes and infer the topological ordering of a DAG. Another line of work employs regression-based methods to detect nonlinearity. For instance, the nonlinear invariant causal prediction (ICP) framework (heinze2018invariant) fits regression models to data collected in different environments, and then tests for differences in residual distributions across environments after interventions to detect associations. gretton2009nonlinear consider the cases where the nonlinear function $f_{i}$ is invertible or where the noise is not additive, and demonstrate that only a PDAG is identifiable in these cases. Their method thereby focuses on identifying non-invertible ANMs in a Markov equivalence class (MEC) through iterative residual independence testing. Improving upon an estimated CPDAG, wang_zhou_nonlinear propose a statistical test that orients the undirected edges by comparing the goodness of fit of models corresponding to the two possible directions. With minimal assumptions on the regression model, this approach offers flexibility in modeling and greater applicability to real-world data.
@@
However, these constraint-based methods also face practical limitations that hinder their widespread adoption. A significant drawback to the iterative testing approach is the computational cost, notably for kernel-based methods. As its runtime scales quadratically with sample size, the KCI test is computationally expensive to employ in constraint-based algorithms, and especially inefficient for learning larger DAGs (strobl2019kci_rcot). Additionally, the performance of the KCI test heavily depends on the chosen kernel function, which is often problem-specific and difficult to tune. The nonlinear ICP requires a sufficient number of observed environments to detect causal relations (rosenfeld2021risks) and thus cannot be applied to data generated under a classical i.i.d. setting. Moreover, the accuracy of regression-based approaches is contingent on the quality of the estimated model, which requires strong domain knowledge for model selection and large sample sizes to accurately approximate the true SEM (shah2020hardness). Model misspecification can lead to violations of key assumptions, ultimately inflating false positive rates in independence tests (li2020nonparametric). More broadly, recent efforts tend to focus on developing independence tests – without fully developing them into scalable, full-fledged causal discovery algorithms (hasan2023survey) – or approximating key statistical quantities to improve test efficiency, with less emphasis on using these metrics to infer causal relations.
@@
Score-based methods for nonlinear learning have also seen significant development in recent years. The algorithm CAM (cam), for instance, assumes an additive noise model with Gaussian errors and maximizes the joint log-likelihood function to learn a DAG. More recently, several works have reformulated the structural learning task as a continuous-optimization problem by devising an algebraic characterization of DAGs (notears). Prominent examples include NOTEARS (zheng2020learning), DAG-GNN (yu2019dag), and DAGMA (bello2023dagmalearningdagsmmatrices), which incorporate deep learning techniques to enhance the flexibility of SEM estimation. Another line of work, exemplified by the SCORE algorithm (rolland2022score), iteratively identifies leaf nodes in a DAG using the Jacobian of the score function. However, score-based methods generally heavily depend on model assumptions, require intensive computational time, and are not guaranteed to find a globally optimal structure. Some of the above deep learning-based methods, in particular, perform poorly when the data is standardized, as standardization erases causal order information from the marginal variance when minimizing the least squares objective function (cd_benchmarks_neurips2021).
@@ 1.2 Contribution of This Work
In lieu of inferring the causal order of nodes, we introduce a local identifiability criterion based on the pairwise additive noise model (PANM). This criterion determines whether a given undirected edge in a partially directed acyclic graph (PDAG) can be correctly oriented. Edges that fulfill this criterion are ensured to be correctly oriented in the large-sample limit, without inducing errors in subsequent orientations. We prove that, at the population level, the algorithm consistently recovers the true DAG from a CPDAG. Leveraging this result, we devise a sequential algorithm to identify undirected edges following the PANM criterion and infer their causal directions. As such, the algorithm effectively learns the DAG at a local scale by evaluating edges individually without necessitating evaluation of all nodes or the entire graph. To determine the orientation of each edge, SNOE employs a likelihood-ratio test that compares the bivariate conditional probability distributions over the sub-DAG on both nodes and their learned parent sets under the competing directions. Contrary to general independence tests, the likelihood ratio test returns a definitive decision regarding the causal direction, thereby bridging the gap between the task of detecting nonlinear conditional independence relations and the structural learning problem. Empirical results further show that the test is robust to violations of model misspecification and yields accurate results across different functional settings.
@@
  • A novel criterion derived from a pairwise additive noise model to determine the identifiability and correct orientation of undirected edges in a PDAG;
  • An algorithm that is guaranteed at the population level to identify the true DAG from its Markov equivalence class by orienting undirected edges, according to their adherence to the PANM criterion, in a sequential manner;
@@
At a conceptual level, the sequential orientation algorithm is the most significant contribution of this work. In essence, our method is rooted in this central idea: Starting from the CPDAG, there exists at least one undirected edge whose orientation can be determined by the PANM criterion at any iteration in our sequential algorithm, leading to the correct recovery of the true DAG. In this process, we check whether the PANM is satisfied for an undirected edge only conditional on the identified parents of the two nodes. This is made possible with a careful design of edge orientations by the PANM criterion and by a subset of the Meek’s rules (meek1995causal_background_knowledge). Although gretton2009nonlinear also orients edges in a CPDAG, they use a greedy search to iteratively identify nodes satisfying a non-invertible ANM by examining all candidate parent sets of a node, i.e., by checking all subsets of its neighbors (connected to the node by an undirected edge). This is clearly different from our sequential orientation procedure that only considers identified parents without the need to examine any subset of the neighbors. The method proposed by wang_zhou_nonlinear ranks the undirected edges in a CPDAG by a goodness-of-fit metric, without establishing its rigorous property in terms of edge orientation or recovery of the true DAG. Moreover, they assume a specific piecewise linear SEM, while we consider a more general nonlinear ANM. resit_anm employs independence tests to identify a sink node in a sequential manner, which is distinct from the above sequential edge orientation methods.
The paper is structured as follows. Section 2 introduces the fundamental concepts and model assumptions. In Section 3, we present the central notion of our work, the sequential edge orientation procedure and the finite-sample version of the algorithm. In Section 4, we discuss the two key components of the edge orientation step: the ranking procedure and the orientation test. Then, we establish the structural learning consistency of the algorithm in Section 5. Section 6 presents the performance of our method against competing methods on simulated data, with a detailed analysis of the algorithm’s intermediate results. We further demonstrate its performance in causal discovery on real-world data in Section 7. Last, we summarize our work and outline directions for future research in Section 8. Proofs and supplemental numerical results are provided in the Appendix.
@@ 2.1 Directed Acyclic Graph
A graph $\mathcal{G}=(V,E)$ consists of a set of vertices $V=[p]:=\{1,\ldots,p\}$ and a set of edges $E\subseteq V\times V$. For a pair of distinct nodes $i,j\in V$, a directed edge $i\rightarrow j\in E$ indicates that node $i$ is a parent of node $j$ and $j$ is a child of $i$. The parent set of node $i$ is denoted by $\text{pa}_{\mathcal{G}}(i)$ and the child set denoted by $\text{ch}_{\mathcal{G}}(i)$. In contrast, there may exist an undirected edge $i-j\in E$ in $\mathcal{G}$, where $i$ is called a neighbor of $j$, i.e. $i\in\text{ne}_{\mathcal{G}}(j)$, and vice versa. A directed acyclic graph (DAG) $\mathcal{G}$ consists only of directed edges and does not admit any directed cycles. A related type of graph is the partially directed acyclic graph (PDAG), which contains both directed and undirected edges but does not contain any cycles in its directed subgraphs.
@@
To infer the structure of the graph $\mathcal{G}$ from the observed data, we require the causal sufficiency and faithfulness assumptions to hold (pearl@2000). Causal sufficiency is satisfied when all common causes of any distinct pair of nodes are observed. In other words, there are no unobserved or latent confounders under this assumption. Suppose $A,B,C\subset[p]$ are disjoint subsets. A DAG $\mathcal{G}$ and the probability distribution $P$ are faithful to one another if the conditional independence relations in $P$ have a one-to-one correspondence with the d-separation relations in $\mathcal{G}$: $A\perp\!\!\!\perp B|C\text{ in }P\Longleftrightarrow C$ d-separates $A,B$ in $\mathcal{G}$. Faithfulness also implies the causal minimality condition. The pair $(\mathcal{G},P)$ satisfies the causal minimality condition if $P$ is not Markov to any proper subgraph of $\mathcal{G}$ over the vertex set $V$ (causation_search_prediction).
@@ 2.2 Markov Equivalence Class
Two DAGs are Markov equivalent if and only if they have identical skeletons and v-structures, which are ordered triplets of nodes $i,j,k$ oriented as $i\rightarrow k\leftarrow j$ with no edge between $i,j$ (verma1990). Markov equivalent DAGs encode the same set of d-separations and form an equivalence class. Without further restrictions on the function classes in the SEM (1), Markov equivalent DAGs cannot be distinguished through observational data; hence we can only learn their equivalence class. The equivalence class is represented by a completed partially directed acyclic graph (CPDAG), a PDAG with specific structural properties (andersson1997characterization). Every directed edge is compelled, or strongly protected, and every undirected edge is reversible in the CPDAG.
The CPDAG $\mathcal{E}$ of a DAG $\mathcal{G}$ is typically obtained by first identifying v-structures in the skeleton, and then applying Meek’s rules, a set of four rules that orient edges based on graphical patterns (meek1995causal_background_knowledge). Meek’s rules identify additional directed edges in the graph without introducing new v-structures. A maximally oriented PDAG is a PDAG for which no edges can be further oriented by Meek’s rules. As an example, a CPDAG $\mathcal{E}$ is a maximally oriented PDAG, or a maximal PDAG for short.
A consistent extension of a PDAG $\mathcal{G}$ is a DAG, $\widetilde{\mathcal{G}}$, obtained by orienting all undirected edges in $\mathcal{G}$ without introducing new v-structures. Therefore, $\widetilde{\mathcal{G}}$ has the same skeleton, the same orientations of all directed edges in $\mathcal{G}$, and the same v-structures as $\mathcal{G}$ (Dor1992ASA). While not all PDAGs can be extended to a DAG, a CPDAG $\mathcal{E}$ is extendable since each DAG in the equivalence class represented by $\mathcal{E}$ is a consistent extension of $\mathcal{E}$. A DAG may be obtained from a CPDAG by iteratively making edge orientations without introducing new v-structures and applying Meek’s rules, while preserving the compelled edges (pmlr-v161-wienobst21a-extendability).